An object is vertically projected at u velocity ..After t time another one is projected at the same velocity. Show the meeting time is (t/2 + u/g) after the first object projected .


1
Expert's answer
2021-05-08T14:39:21-0400

"t_{up}=\frac ug,~s_{up}=\frac{u^2}{2g},"

"s=s_{up}-s'=\frac{u^2}{2g}-u(\frac ug-t)+\frac{gt^2}{2}=ut-\frac{u^2}{2g}+\frac{gt^2}{2},"

"v'=u-g(\frac ug-t)=gt,"

"gtt'-\frac{gt'^2}{2}=s-\frac{gt'^2}{2},"

"t'=\frac{s}{gt}=\frac ug-\frac t2,"

"T=t+t'=\frac t2+\frac ug."